# QuantFace: Low-Bit Post-Training Quantization for One-Step Diffusion Face Restoration

### v1

@@ 3.2 Rotation-Scaling Channel Balancing
To minimize the additional computational overhead, inspired by the offline transformation strategies from [42, 45], we integrate most of the scaling factors and rotation matrices into model weights, thereby minimizing the quantization overhead. During inference, we only perform online RHT for input activations of convolution layers, incurring only a 2.15% additional computational overhead.
3.3 Quantization-Distillation Low-Rank Adaptation
@@
Under extremely low bit-width configurations, generative models face severe difficulty in producing viable outputs. Incorporating a high-precision branch to mitigate quantization errors is a widely adopted strategy [15, 21]. SVDQuant [15] shows that initializing these branches via singular value decomposition (SVD) [46] has proven to be particularly effective. Inspired by SVDQuant, we use a low-rank branch initialized by SVD to reduce the quantization error. This can be formulated as follows [15]:
$XW=XL_{1}L_{2}+XR\approx XL_{1}L_{2}+Q(X)\,Q(R),\quad R=W-L_{1}L_{2},$ (9)
@@
where the SVD of $W=U\Sigma V$, the $r$ rank optimal solution is $L_{1}=U\Sigma_{:,:r},L_{2}=V_{:r,:}$. However, most update rules for low-rank branches used in previous works are heuristic [15, 47], and do not guarantee convergence. Meanwhile, existing approaches predominantly concentrate on minimizing per-layer quantization error and often neglect the final results. Moreover, previous studies [48, 49] find that minimizing per-layer reconstruction error does not necessarily optimize the generation quality. We need a more suitable update scheme for the low-rank branches.
@@
To solve this problem, a straightforward alternative is Low-Rank Adaptation (LoRA) [50], a Parameter-Efficient Fine-Tuning method, to align the quantized model with the FP model. However, although this approach is effective, the resulting image quality remains unsatisfactory. Inspired by FBquant [51], we confirmed that training the SVD-initialized low-rank branch would regularize the model to not deviate too far from the original one. But meanwhile, this strong constraint impedes distillation from the FP model to the quantized model. To remedy this problem, we introduce additional degrees of freedom into optimization. Therefore, we are able to conduct effective model fine-tuning.
@@
Specifically, we propose Quantization-Distillation Low-Rank Adaptation (QD-LoRA). We train two low-rank branches simultaneously, while maintaining an identical number of parameters with SVDQuant [15]. The first branch is initialized via SVD and constrained by the backbone weights during training. It is designed to prevent excessive deviation from the original parameters. The second branch utilizes standard LoRA initialization and is dedicated to distilling knowledge from the FP model to the quantized one. The inference procedure is formulated as follows:
$XW=XL_{1}L_{2}+XR\approx XAB+XL_{1}L_{2}+Q(X)\,Q(R),\quad R=W-L_{1}L_{2},$ (10)
@@
where we use a random Gaussian initialization for $A$ and zero for $B$. After training, the two low-rank branches can be merged together. That means the auxiliary low-rank branch does not introduce any additional computational overhead during inference, compared with SVDQuant [15].
@@
In practice, we find that QD-LoRA achieves superior face restoration performance compared to optimizing a single low-rank branch while maintaining an identical number of parameters.
@@ 3.4 Adaptive Bit-width Allocation
Another challenge in low-bit quantization is the suboptimal allocation of limited resources. Previous works [20] notice that different layers exhibit varying sensitivities to quantization. We measure the change in FID [52] on FFHQ [53] by quantizing the activations of each layer individually, and we observe analogous findings in activation quantization.
As illustrated in Fig. 5, within the OSDFace UNet, key convolution layers and residual connections have an outsized impact on the final image quality. After identifying the quantization bottlenecks, we can then devise a mixed-precision scheme. We assign a greater bit-width to those sensitive layers while keeping the resource budget unchanged.
Enlightened by prior works [54, 20, 55], we model the bit-width allocation as an integer programming problem, which is called adaptive bit-width allocation. The integer programming problem can be solved efficiently, compared with other methods like genetic algorithms [56].
When designing our integer programming objective, we focus on both reconstruction error and perception metric. Prior works [44, 20] observe that simply allocating higher bit-widths to layers with large reconstruction error (i.e., MSE) does not guarantee better visual performance. Consequently, we employ a heuristic approach to assign perceptual error weight to each activation layer. Specifically, we first quantize the activation of every single layer to 2 bits, while keeping the remaining activation values in full precision. We then progressively increase the bit-width of the quantized activation until the FID difference compared to the FP model is within $\epsilon$, or until the maximum bit-width $B_{max}$ is reached. Then, for each layer $i$, we extract the corresponding bit-width $B_{i}$ and assign the perceptual error weight as $w_{i}=2^{B_{i}}-1$. In practice, we set $\epsilon=0.02,B_{max}=8$.
@@
The Detailed integer programming workflow can be formulated as follows [54]:
@@
[…]
@@
Table 1: Quantitative comparison on the synthetic dataset CelebA-Test [57]. The best and second best results are colored with red and blue.
[…]
32/32 OSDFace [12] 0.6910 0.1773 0.3365 0.5423 75.64 3.884 60.0708 5.287 45.41 17.06
MaxMin [38] 0.1781 0.4022 0.6082 0.1406 12.95 11.512 79.7129 11.214 234.94 216.67
Q-Diffusion [17] 0.3679 0.2853 0.4784 0.1770 24.23 6.821 73.5904 7.840 81.43 75.69
EfficientDM [21] 0.4571 0.2493 0.4128 0.3820 53.75 7.269 63.4590 5.800 63.47 34.77
PassionSR [14] 0.6309 0.1762 0.3368 0.4731 72.74 4.125 61.5067 5.633 43.08 23.00
SVDQuant [15] 0.6566 0.1774 0.3397 0.4818 73.59 4.021 60.7007 5.501 44.57 22.44
6/6 QuantFace 0.6539 0.1740 0.3339 0.4893 74.43 4.045 60.1039 5.363 41.84 19.23
MaxMin [38] 0.1694 0.4144 0.6354 0.1472 12.93 12.864 80.8289 14.038 261.82 251.00
Q-Diffusion [17] 0.2399 0.3541 0.5505 0.1141 16.45 9.091 77.0839 9.125 136.06 118.48
EfficientDM [21] 0.4981 0.2475 0.4113 0.4119 55.96 7.212 64.8563 6.096 60.11 33.96
PassionSR [14] 0.6391 0.1803 0.3401 0.4719 72.38 4.326 63.3448 5.737 45.73 26.01
SVDQuant [15] 0.6912 0.1884 0.3639 0.4732 71.79 3.762 63.3879 6.043 50.51 33.18
4/6 QuantFace 0.6510 0.1770 0.3442 0.4612 72.83 4.127 61.7154 5.658 42.18 23.36
MaxMin [38] 0.2045 0.4466 0.6867 0.2137 13.45 18.011 82.1358 16.783 315.62 311.01
Q-Diffusion [17] 0.1747 0.4258 0.6660 0.1838 13.26 16.620 82.0933 17.066 295.58 289.39
EfficientDM [21] 0.1992 0.4326 0.6368 0.1692 13.05 15.275 81.6171 15.463 265.51 254.24
PassionSR [14] 0.3248 0.3426 0.5603 0.1398 15.66 6.821 82.4272 18.233 250.54 279.35
SVDQuant [15] 0.5192 0.2563 0.4143 0.3572 58.90 5.709 76.3504 10.387 84.91 72.36
4/4 QuantFace 0.6045 0.1867 0.3485 0.4255 71.49 4.813 67.5026 5.951 55.16 32.52
4/3.98 QuantFace-MP 0.5998 0.1830 0.3383 0.4491 73.00 4.708 65.7544 5.860 51.02 27.72
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 2862 LQ - 2862 OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours) Bits (W/A) W32A32 W4A6 W4A6 W4A6 W4A6 W4A6 W4A6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 0254 LQ - 0254 OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours) Bits (W/A) W32A32 W4A4 W4A4 W4A4 W4A4 W4A4 W4A3.98
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 0120 LQ - 0120 OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours) Bits (W/A) W32A32 W4A4 W4A4 W4A4 W4A4 W4A4 W4A3.98
@@ 4.1 Experiment Settings
We randomly select 512 high-quality face images from FFHQ [53] and resized them to 512$\times$512 pixels. We obtained our synthetic training data using a dual-stage degradation model which is consistent with WaveFace [58]. The synthetic data collected in this manner closely aligns with the distribution of low-quality face images observed in real-world scenarios. we evaluate our quantized model on both synthetic and real-world datasets. The synthetic dataset is CelebA-Test [57] from DAEFR [27] and the real-world datasets are Wider-Test [59], WebPhoto-Test [60], and LFW-Test [61], which are consistent with OSDFace [12].
@@
For the Synthetic Datasets, LPIPS [62] and DISTs [63] are adopted as reference-based perceptual measures, together with CLIPIQA [64], MANIQA [65], NIQE [66], and MUSIQ [67] as non-reference metrics. FID [52] with both FFHQ [53] and CelebA-Test HQ [57] are used to evaluate the distribution similarity between the real faces and generated ones. Moreover, following the previous face restoration works [24, 59, 27], we also evaluate the embedding angle of ArcFace [68] called ‘Deg.’, and the landmark distance named ‘LMD’. For the Real-world Datasets, CLIPIQA [64], MANIQA [65], MUSIQ [67], NIQE [66], and FID [52] with FFHQ [53] are employed.
@@
Learning rate for QuantFace is set to $10^{-5}$ using the Adam [69] optimizer. Only the low-rank branches of each layer are updated during training, and the LoRA rank for both branches is 8. The loss function jointly integrates a perceptual loss in the pixel domain with a numerical reconstruction loss, which is formulated as:
$\mathcal{L}$ $=\lambda_{1}\mathcal{L}_{\text{lpips}}(I_{q},I_{fp})+\lambda_{2}\mathcal{L}_{mse}(I_{q},I_{fp}),\vspace{-3mm}$ (12)
here, we set $\lambda_{1}=0.5,\lambda_{2}=0.5$ empirically. Experiments are conducted on a single NVIDIA A800 GPU, consuming 35.7 GB of GPU memory and 2.6 hours of GPU time.
Compard Methods.
We compare QuantFace with several representative quantization methods: MaxMin [38], Q-Diffusion [17], EfficientDM [21], PassionSR [14], and SVDQuant [15]. Q-Diffusion [17], EfficientDM [21], and SVDQuant [15] are quantization methods for multi-step diffusion models. PassionSR [14] is a quantization method designed for one-step diffusion models.
Table 2: Params, Ops, and compression ratio of various settings. Ops are computed with input size 512$\times$512.
@@
OSDFace [12] W32A32 978 ($\downarrow$0%) 4,264 ($\downarrow$0%)
W6A6 209 ($\downarrow$78.61%) 995 ($\downarrow$76.66%)
W4A6 148 ($\downarrow$84.85%) 848 ($\downarrow$80.10%)
QuantFace W4A4 148 ($\downarrow$84.85%) 728 ($\downarrow$82.91%)
@@
To evaluate computational cost, following previous work [70], we measure the total model size (Params / M) and the number of operations (Ops / G). The results are summarized in Tab. 2. Under the 4-bit setting, our QuantFace achieves approximately 84.85% compression ratio for parameters and 82.91% compression ratio for computation.
@@
[…]

### v2

@@ 3.2 Rotation-Scaling Channel Balancing
To minimize the additional computational overhead, inspired by the offline transformation strategies from [quarot, Xu2025MambaQuantQT], we integrate most of the scaling factors and rotation matrices into model weights, thereby minimizing the quantization overhead. During inference, we only perform online RHT for input activations of convolution layers.
3.3 Quantization-Distillation LoRA
@@
Under extremely low bit-width configurations, face restoration models struggle to generate fine facial detail features. Incorporating a high-precision branch to mitigate quantization errors is a widely adopted strategy [li2025svdquant, He2023EfficientDM]. SVDQuant [li2025svdquant] shows that initializing these branches via singular value decomposition (SVD) [svd] has proven to be particularly effective. Inspired by SVDQuant, we use a low-rank branch initialized by SVD to reduce the quantization error. This can be formulated as follows [li2025svdquant]:
$XW=XL_{1}L_{2}+XR\approx XL_{1}L_{2}+Q(X)\,Q(R),$ (9)
@@
where the SVD of $W=U\Sigma V$, the $r$ rank optimal solution is $L_{1}=U\Sigma_{:,:r},L_{2}=V_{:r,:}$, and $R=W-L_{1}L_{2}$. However, most update rules for low-rank branches used in previous works are heuristic [li2025svdquant, guo2023lq], and do not guarantee convergence. Meanwhile, existing approaches predominantly concentrate on minimizing per-layer quantization error and often neglect the final results. Moreover, previous studies [zhang2024qera, deng2023mixed] find that minimizing per-layer reconstruction error does not necessarily optimize the generation quality. We need a more suitable update scheme for the low-rank branches.
@@
To solve this problem, a straightforward alternative is Low-Rank Adaptation (LoRA) [hu2022lora], a Parameter-Efficient Fine-Tuning method, to align the quantized model with the FP model. However, although this approach is effective, the resulting image quality remains unsatisfactory. Inspired by Fbquant [liu2025fbquant], we confirmed that training the SVD‑initialized low‑rank branch would regularize the model to not deviate too far from the original one. Meanwhile, this strong constraint impedes distillation from the FP model to the quantized model. To remedy this problem, we introduce additional degrees of freedom into optimization. Therefore, we can conduct effective model fine-tuning.
@@
Specifically, we propose Quantization-Distillation Low-Rank Adaptation (QD-LoRA). We train two low-rank branches simultaneously, while maintaining an identical number of parameters with SVDQuant [li2025svdquant]. The first branch is initialized via SVD and constrained by the backbone weights during training. It is designed to prevent excessive deviation from the original parameters. The second branch utilizes standard LoRA initialization and is dedicated to distilling knowledge from the FP model to the quantized one. The inference procedure is formulated as follows:
$XW$ $=XL_{1}L_{2}+XR$ (10)
$\approx XAB+XL_{1}L_{2}+Q(X)\,Q(R),$
@@
where we use a random Gaussian initialization for $A$ and zero for $B$ and $R=W-L_{1}L_{2}$. After training, the two low-rank branches can be merged. The auxiliary low-rank branch does not introduce any additional computational overhead during inference compared with SVDQuant [li2025svdquant].
@@
In practice, we find that QD-LoRA achieves superior face restoration performance compared to optimizing a single low-rank branch with an identical number of parameters.
[…]
32/32 OSDFace [osdface] 0.6910 0.1773 0.3365 0.5423 75.64 3.884 60.0708 5.287 45.41 17.06
MaxMin [jacob2017quantizationtrainingneuralnetworks] 0.1781 0.4022 0.6082 0.1406 12.95 11.512 79.7129 11.214 234.94 216.67
Q-Diffusion [qdiffusion] 0.3679 0.2853 0.4784 0.1770 24.23 6.821 73.5904 7.840 81.43 75.69
EfficientDM [He2023EfficientDM] 0.4571 0.2493 0.4128 0.3820 53.75 7.269 63.4590 5.800 63.47 34.77
PassionSR [zhu2024passionsr] 0.6309 0.1762 0.3368 0.4731 72.74 4.125 61.5067 5.633 43.08 23.00
SVDQuant [li2025svdquant] 0.6566 0.1774 0.3397 0.4818 73.59 4.021 60.7007 5.501 44.57 22.44
6/6 QuantFace 0.6670 0.1783 0.3345 0.5122 74.89 3.759 60.5890 5.413 42.99 19.02
MaxMin [jacob2017quantizationtrainingneuralnetworks] 0.1694 0.4144 0.6354 0.1472 12.93 12.864 80.8289 14.038 261.82 251.00
Q-Diffusion [qdiffusion] 0.2399 0.3541 0.5505 0.1141 16.45 9.091 77.0839 9.125 136.06 118.48
EfficientDM [He2023EfficientDM] 0.4981 0.2475 0.4113 0.4119 55.96 7.212 64.8563 6.096 60.11 33.96
PassionSR [zhu2024passionsr] 0.6391 0.1803 0.3401 0.4719 72.38 4.326 63.3448 5.737 45.73 26.01
SVDQuant [li2025svdquant] 0.6912 0.1884 0.3639 0.4732 71.79 3.762 63.3879 6.043 50.51 33.18
4/6 QuantFace 0.6660 0.1804 0.3329 0.5317 75.33 4.128 62.0754 5.457 49.59 19.51
MaxMin [jacob2017quantizationtrainingneuralnetworks] 0.2045 0.4466 0.6867 0.2137 13.45 18.011 82.1358 16.783 315.62 311.01
Q-Diffusion [qdiffusion] 0.1747 0.4258 0.6660 0.1838 13.26 16.620 82.0933 17.066 295.58 289.39
EfficientDM [He2023EfficientDM] 0.1992 0.4326 0.6368 0.1692 13.05 15.275 81.6171 15.463 265.51 254.24
PassionSR [zhu2024passionsr] 0.3248 0.3426 0.5603 0.1398 15.66 6.821 82.4272 18.233 250.54 279.35
SVDQuant [li2025svdquant] 0.5192 0.2563 0.4143 0.3572 58.90 5.709 76.3504 10.387 84.91 72.36
4/4 QuantFace 0.6916 0.1914 0.3709 0.5222 73.69 4.072 69.2888 6.195 58.34 31.17
4/3.97 QuantFace-MP 0.6984 0.1869 0.3518 0.5560 75.17 4.116 65.6640 5.927 53.18 25.08
Table 1: Quantitative comparison on the synthetic dataset CelebA-Test [CelebA-Test]. The best and second best results are colored with red and blue.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 2862 LQ - 2862 OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours) Bits (W/A) W32A32 W4A6 W4A6 W4A6 W4A6 W4A6 W4A6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 0254 LQ - 0254 OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours) Bits (W/A) W32A32 W4A4 W4A4 W4A4 W4A4 W4A4 W4A3.97
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HQ - 0969 LQ - 0969 OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours) Bits (W/A) W32A32 W4A4 W4A4 W4A4 W4A4 W4A4 W4A3.97
@@ 3.4 Adaptive Bit-width Allocation
Another challenge in low-bit quantization is the suboptimal allocation of limited resources. Previous works [sui2024bitsfusion] notice that different layers exhibit varying sensitivities to quantization. We measure the change in FID [fid] on FFHQ [ffhq] by quantizing the activations of each layer individually, and we observe analogous findings in activation quantization.
As illustrated in Fig. 5, within the OSDFace UNet, key convolution layers and residual connections have an outsized impact on the final image quality. After identifying the quantization bottlenecks, we can then devise a mixed‑precision scheme. Enlightened by prior works [zhao2024mixdq, sui2024bitsfusion, ma2023ompq], we model the bit-width allocation as an integer programming problem, which is called adaptive bit-width allocation.
When designing our integer programming objective, we focus on both reconstruction error and perception metric. Prior works [vidit, sui2024bitsfusion] observe that simply allocating higher bit‑widths to layers with large reconstruction error (i.e., MSE) does not guarantee better visual performance. Consequently, we employ a heuristic approach to assign perceptual error weight to each activation layer. Specifically, we first quantize the activation of every single layer to 4-bit. We then progressively increase the bit-width of the quantized activation until the FID difference compared to the FP model is within $\epsilon=0.02$, or until the maximum bit-width $B_{max}=8$ is reached. Then, for each layer $i$, we extract the corresponding bit-width $B_{i}$ and assign the weight as $w_{i}=2^{B_{i}}-1$.
@@
The Detailed integer programming workflow can be formulated as follows [zhao2024mixdq]:
@@
$\text{s.t.}\quad\sum_{b=3,4,5,6}a_{i,b}=1,\quad\sum_{i=1}^{N}\sum_{b=3,4,5,6}a_{i,b}\cdot\mathcal{M}_{i,b}\leq\mathcal{B},$
@@ 4.1 Experiment Settings
We randomly select 1000 high-quality face images from FFHQ [ffhq] and resized them to $\times512$ pixels. We obtained our synthetic training data using a dual-stage degradation model, which is consistent with WaveFace [waveface]. We evaluate our quantized model on both synthetic and real‑world datasets. The synthetic dataset is CelebA-Test [CelebA-Test] from DAEFR [daefr], and the real-world datasets are Wider-Test [wider-test], WebPhoto-Test [webphoto-test], and LFW-Test [lfw-test], which are consistent with OSDFace [osdface].
@@
For the Synthetic Datasets, LPIPS [lpips] and DISTS [dists] are adopted as reference-based perceptual measures, together with CLIPIQA [clipiqa], MANIQA [maniqa], NIQE [niqe], and MUSIQ [musiq] as non-reference metrics. FID [fid] with both FFHQ [ffhq] and CelebA-Test HQ [CelebA-Test] are used to evaluate the distribution similarity between the real faces and generated ones. Moreover, following the previous face restoration works [vqfr, wider-test, daefr], we also evaluate the embedding angle of ArcFace [arcface] called ‘Deg.’, and the landmark distance named ‘LMD’. For the Real-world Datasets, CLIPIQA [clipiqa], MANIQA [maniqa], MUSIQ [musiq], NIQE [niqe], and FID [fid] with FFHQ [ffhq] are employed.
@@
Learning rate for QuantFace is set to $10^{-5}$ using the Adam [adam] optimizer. Only the low-rank branches of each layer are updated during training, and the LoRA rank for both branches is 16. The overall loss function for the generator is defined as
$\mathcal{L}_{\text{gen}}$ $=\lambda_{\text{dis}}\cdot\mathcal{L}_{\mathcal{G}}(z_{q},z_{fp})+\lambda_{\text{ID}}\cdot\mathcal{L}_{\text{ID}}(I_{q},I_{fp})$ (12)
$+\lambda_{\text{per}}\cdot\mathcal{L}_{\text{LPIPS}}(I_{q},I_{fp})+\lambda_{mse}\cdot\operatorname{MSE}(I_{q},I_{fp}).$
To keep the face identity, we utilize a pretrained ArcFace model [arcface] to encode both the quantized and FP faces into identity embeddings, where
$\mathcal{L}_{\text{ID}}=1-\cos\left(\mathcal{F}(I_{q}),\mathcal{F}(I_{fp})\right).$ (13)
Furthermore, to enhance the quality of quantized faces, we adopt the same GAN training strategy as that employed in OSDFace [osdface], where
$\mathcal{L}_{\mathcal{G}}$ $=-\mathbb{E}_{t}\left[\log\mathcal{D}_{\theta}\left(F(z_{q},t)\right)\right],$ (14)
$\mathcal{L}_{\mathcal{D}}$ $=-\mathbb{E}_{t}\left[\log\left(1-\mathcal{D}_{\theta}\left(F(z_{q},t)\right)\right)\right]$
$-\mathbb{E}_{t}\left[\log\mathcal{D}_{\theta}\left(F(z_{fp},t)\right)\right],$
Experiments are conducted on a single NVIDIA RTX A6000 GPU, consuming 29 GB of GPU memory and 6 hours of GPU time for training.
Compare Methods.
We compare QuantFace with several representative quantization methods: MaxMin [jacob2017quantizationtrainingneuralnetworks], Q-Diffusion [qdiffusion], EfficientDM [He2023EfficientDM], PassionSR [zhu2024passionsr], and SVDQuant [li2025svdquant]. Q-Diffusion [qdiffusion], EfficientDM [He2023EfficientDM], and SVDQuant [li2025svdquant] are quantization methods for multi-step diffusion models. PassionSR [zhu2024passionsr] is a quantization method designed for one-step diffusion models.
@@
OSDFace [osdface] W32A32 866 ($\downarrow$0%) 678 ($\downarrow$0%)
W6A6 185 ($\downarrow$78.61%) 158 ($\downarrow$76.66%)
W4A6 131 ($\downarrow$84.85%) 135 ($\downarrow$80.10%)
QuantFace W4A4 131 ($\downarrow$84.85%) 116 ($\downarrow$82.91%)
Table 2: Params, Ops, and compression ratio (UNet only) of various settings. Ops are computed with input size $\times512$.
@@
To evaluate computational cost, following previous work [qin2023quantsr], we measure the total model size (Params / M) and the number of operations (Ops / G). The results are summarized in Tab. 2. Under the 4-bit setting, our QuantFace achieves approximately 84.85% compression ratio for parameters and 82.91% compression ratio for computation, compared with the original UNet.
[…]
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Wider 0022 OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption WebPhoto 00038 OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LFW E. Nordegren OSDFace [osdface] MinMax [jacob2017quantizationtrainingneuralnetworks] Q-Diffusion [qdiffusion] EfficientDM [He2023EfficientDM] PassionSR [zhu2024passionsr] SVDQuant [li2025svdquant] QuantFace (ours)
Figure 7: Visual performance comparison on real-world datasets. Please zoom in for a better view. Our QuantFace is under bit-width W4A3.97, and other methods are under W4A4.